A walkway forms one diagonal of a square playground. The walkway is 14 m long. How long is a side of the​ playground?
Each side of the playground is(Answer) m.
​(Round to the nearest meter as​ needed.)

Answer by Cromlix(4381)

Hi there,
Make the sides = 'x'
By Pythagoras theorem
x^2 + x^2 = 14^2
2x^2 = 196
x^2 = 196/2
x^2 = 98
x = √98
x = 9.9 meters.
The side = 10 meters.
Hope this helps :-)